# Unveiling the Role of Learning Rate Schedules via Functional Scaling Laws

### v1

@@ C.3 The g Functions
$\int_{0}^{t}G_{a}(t-\tau)G_{b}(\tau)\mathop{}\!\mathrm{d}\tau\eqsim G_{a\wedge b}(t).$
C.4 The Case of Top-M Features
In this section, we prove the functional scaling law for the case of top-$M$ features. Recall the choice of top-$M$ features, where we choose $\mathbf{W}=[\mathbf{I}_{M},\bm{0}_{M\times(N-M)}]$. In this case we can simplify the SDE and the expression of the $\mathcal{E}_{t}$. Then we will estimate each term of the Volterra equation to complete the proof.
Lemma C.11 (The intrinsic-time SDE of top-$M$ case).
Let the projector $\mathbf{W}=[\mathbf{I}_{M},\bm{0}_{M\times(N-M)}]$. Then, we have
$\mathop{}\!\mathrm{d}\mathbf{u}_{t}=-\mathbf{H}_{0:M}\mathbf{u}_{t}\mathop{}\!\mathrm{d}t+\sqrt{\gamma_{t}(c_{t}\mathcal{E}_{t}+\sigma^{2})}\sqrt{\mathbf{H}_{0:M}}\mathop{}\!\mathrm{d}\mathbf{B}_{t}.$ (19)
where $\mathbf{H}_{0:M}=\operatorname{diag}\{\underbrace{\lambda_{1},...,\lambda_{M}}_{M},\underbrace{0,...,0}_{N-M}\}$.
Proof.
Recall that from Eq. (15),
$\mathop{}\!\mathrm{d}\mathbf{u}_{t}=-\mathbf{W}^{\top}\mathbf{W}\mathbf{H}\mathbf{u}_{t}\mathop{}\!\mathrm{d}t+\sqrt{\gamma_{t}(c_{t}\mathcal{E}_{t}+\sigma^{2})}\sqrt{\mathbf{W}^{\top}\mathbf{W}\mathbf{H}\mathbf{W}^{\top}\mathbf{W}}\mathop{}\!\mathrm{d}\mathbf{B}_{t}$
Notice that
$\mathbf{W}^{\top}\mathbf{W}=\operatorname{diag}\{\underbrace{1,...,1}_{M},\underbrace{0,...,0}_{N-M}\},$
and we have
$\mathbf{W}^{\top}\mathbf{W}\mathbf{H}=\mathbf{H}_{0:M},\qquad\mathbf{W}^{\top}\mathbf{W}\mathbf{H}\mathbf{W}^{\top}\mathbf{W}=\mathbf{H}_{0:M}.$
Combine these results and we have
$\mathop{}\!\mathrm{d}\mathbf{u}_{t}=-\mathbf{H}_{0:M}\mathbf{u}_{t}\mathop{}\!\mathrm{d}t+\sqrt{\gamma_{t}(c_{t}\mathcal{E}_{t}+\sigma^{2})}\sqrt{\mathbf{H}_{0:M}}\mathop{}\!\mathrm{d}\mathbf{B}_{t}.$
We have the following recursion for $\mathcal{E}_{t}$.
Lemma C.12 (Volterra equation of the top-$M$ case).
We have
$\mathbb{E}[\mathcal{E}_{t}]\eqsim g_{s}(t)+\int_{0}^{t}\mathcal{K}(t-\tau)\gamma_{\tau}\left(c_{\tau}\mathbb{E}[\mathcal{E}_{\tau}]+\sigma^{2}\right)\mathop{}\!\mathrm{d}\tau+M^{-s\beta}.$
Proof.
The proof is mainly based on the Lemma C.11. Similarly, we decompose the SDE into its individual components, solve the resulting equations. By addressing each component individually, we can derive precise solutions and expectations for $(\mathbf{u}_{t}^{(j)})^{2}$. For $1\leqslant j\leqslant M$, we have
$\mathbb{E}[(\mathbf{u}_{t}^{(j)})^{2}]=e^{-2\lambda_{j}t}(\mathbf{v}_{0}^{(j)})^{2}+\lambda_{j}\int_{0}^{t}\gamma_{\tau}e^{-2\lambda_{j}(t-\tau)}(c_{\tau}\mathbb{E}[\mathcal{E}_{\tau}]+\sigma^{2})\mathop{}\!\mathrm{d}\tau,$
We then calculate the risk:
$2\mathbb{E}[\mathcal{E}_{t}]$ $=\sum_{j=1}^{M}\lambda_{j}e^{-2\lambda_{j}t}(\mathbf{v}_{0}^{(j)})^{2}+\sum_{j=1}^{M}\lambda_{j}^{2}\int_{0}^{t}\gamma_{\tau}e^{-2\lambda_{j}(t-\tau)}(c_{\tau}\mathbb{E}[\mathcal{E}(\mathbf{v}_{\tau})]+\sigma^{2})\mathop{}\!\mathrm{d}\tau+\sum_{j=M+1}^{\infty}\lambda_{j}(\mathbf{v}_{0}^{(j)})^{2}$
$=\underbrace{\sum_{j=1}^{M}\lambda_{j}e^{-2\lambda_{j}t}(\mathbf{v}_{0}^{(j)})^{2}}_{I_{1}}+\underbrace{\sum_{j=1}^{M}\lambda_{j}^{2}\int_{0}^{t}\gamma_{\tau}e^{-2\lambda_{j}(t-\tau)}(c_{\tau}\mathbb{E}[\mathcal{E}(\mathbf{v}_{\tau})])\mathop{}\!\mathrm{d}\tau}_{I_{2}}+\underbrace{\sum_{j=1}^{M}\lambda_{j}^{2}\int_{0}^{t}\gamma_{\tau}e^{-2\lambda_{j}(t-\tau)}(\sigma^{2})\mathop{}\!\mathrm{d}\tau}_{I_{3}}$
$+\underbrace{\sum_{j=M+1}^{\infty}\lambda_{j}(\mathbf{v}_{0}^{(j)})^{2}}_{I_{4}}.$
We next shall bound the four terms separately.
  • The $I_{1}$ term. We can substitute $\lambda_{j}$ by $j^{-\beta}$ and approximate the sum by integral
    $I_{1}$ $\eqsim\sum_{j=1}^{M}j^{-s\beta-1}e^{-2\lambda_{j}t}\eqsim\int_{1}^{M}x^{-s\beta-1}e^{-2x^{-\beta}t}\mathop{}\!\mathrm{d}x=\frac{1}{\beta}\int_{M^{-\beta}}^{1}u^{s-1}e^{-2ut}\mathop{}\!\mathrm{d}u\eqsim g_{s}(t).$
  • The $I_{2}$ term. Similarly, we substitute $\lambda_{j}$ by $j^{-\beta}$ and approximate the sum by integral
    $I_{2}$ $=\int_{0}^{t}\gamma_{\tau}c_{\tau}\mathbb{E}[\mathcal{E}_{\tau}]\left(\sum_{j=1}^{M}\lambda_{j}^{2}e^{-2\lambda_{j}(t-\tau)}\mathop{}\!\mathrm{d}\tau\right)\eqsim\int_{0}^{t}\gamma_{\tau}c_{\tau}\mathbb{E}[\mathcal{E}_{\tau}]\left(\int_{1}^{M}x^{2\beta}(e^{-2x^{-\beta}(t-\tau)})dx\right)$
    $=\frac{1}{\beta}\int_{0}^{t}\gamma_{\tau}c_{\tau}\mathbb{E}[\mathcal{E}_{\tau}]\left(\int_{M^{-\beta}}^{1}u^{1-1/\beta}e^{-2u(t-\tau)}du\right)\eqsim\int_{0}^{t}\mathcal{K}(t-\tau)\gamma_{\tau}c_{\tau}\mathbb{E}[\mathcal{E}_{\tau}]\mathop{}\!\mathrm{d}\tau.$ (20)
  • The $I_{3}$ term. Similarly, we substitude $\lambda_{j}$ by $j^{-\beta}$ and approximate the sum by integral
    $I_{3}$ $=\sigma^{2}\int_{0}^{t}\gamma_{\tau}\left(\sum_{j=1}^{M}\lambda_{j}^{2}e^{-2\lambda_{j}(t-\tau)}\mathop{}\!\mathrm{d}\tau\right)\eqsim\sigma^{2}\int_{0}^{t}\gamma_{\tau}\left(\int_{1}^{M}x^{2\beta}(e^{-2x^{-\beta}(t-\tau)})dx\right)$
    $=\frac{1}{\beta}\sigma^{2}\int_{0}^{t}\gamma_{\tau}\left(\int_{M^{-\beta}}^{1}u^{1-1/\beta}e^{-2u(t-\tau)}du\right)\eqsim\int_{0}^{t}\mathcal{K}(t-\tau)\gamma_{\tau}\sigma^{2}\mathop{}\!\mathrm{d}\tau.$
  • The $I_{4}$ term. By approximating the sum with the integral, we have
    $I_{4}=\sum_{j=M+1}^{\infty}j^{-1-s\beta}\eqsim\int_{M}^{\infty}x^{-1-s\beta}\mathop{}\!\mathrm{d}x=M^{-s\beta}.$
Combining them together completes the proof. ∎
Proposition C.13 (Functional Scaling Law of the top-$M$ case).
Suppose that $\gamma_{\max}:=\sup_{t}\gamma_{t}$ is sufficiently small. Then, for $s\in(0,2-1/\beta]$ we have
$\mathbb{E}[\mathcal{E}_{t}]\eqsim g_{s}(t)+\int_{0}^{t}\mathcal{K}(t-\tau)\left(\sigma^{2}+g_{s}(\tau)\right)\gamma_{\tau}\mathop{}\!\mathrm{d}\tau+M^{-s\beta}.$
Proof.
The case of $M^{\beta}\geqslant t$: The condition of Lemma C.9 holds.
From Lemma C.12 we have
@@ C.4 The Case of Top-M Features
$\mathbb{E}[\mathcal{E}_{t}]\eqsim g_{s}(t)+M^{-s\beta}+\int_{0}^{t}\mathcal{K}(t-\tau)c_{\tau}\mathbb{E}[\mathcal{E}_{\tau}]\gamma_{\tau}\mathop{}\!\mathrm{d}\tau+\sigma^{2}\int_{0}^{t}\mathcal{K}(t-\tau)c_{\tau}\gamma_{\tau}\mathop{}\!\mathrm{d}\tau.$
First we define $B(t)$ as
$B(t)$ $=\int_{0}^{t}\mathcal{K}(t-\tau)\,c_{\tau}\mathbb{E}[\mathcal{E}_{\tau}]\gamma_{\tau}\mathop{}\!\mathrm{d}\tau$
$\eqsim\int_{0}^{t}\mathcal{K}(t-\tau)\,c_{\tau}\gamma_{\tau}\left(g_{s}(\tau)+M^{-s\beta}+\sigma^{2}\int_{0}^{\tau}\mathcal{K}(\tau-r)c_{r}\gamma_{r}\mathop{}\!\mathrm{d}r\right)\mathop{}\!\mathrm{d}\tau$
$\qquad+\int_{0}^{t}\mathcal{K}(t-\tau)\,c_{\tau}\gamma_{\tau}\int_{0}^{\tau}\mathcal{K}(\tau-r)c_{r}\mathbb{E}[\mathcal{E}_{r}]\gamma_{r}\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}\tau.$ (21)
Then we calculate the last term. By Fubini’s theorem,
$\int_{0}^{t}\mathcal{K}(t-\tau)\,c_{\tau}\gamma_{\tau}\int_{0}^{\tau}\mathcal{K}(\tau-r)c_{r}\mathbb{E}[\mathcal{E}_{r}]\gamma_{r}\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}\tau$
$=\int_{0}^{t}c_{r}\mathbb{E}[\mathcal{E}_{r}]\gamma_{r}\int_{r}^{t}\mathcal{K}(t-\tau)\mathcal{K}(\tau-r)\gamma_{\tau}\mathop{}\!\mathrm{d}\tau\mathop{}\!\mathrm{d}r$
$\leqslant\int_{0}^{t}c_{r}\mathbb{E}[\mathcal{E}_{r}]\gamma_{r}\int_{r}^{t}\mathcal{K}(t-\tau)\mathcal{K}(\tau-r)\gamma_{\max}\mathop{}\!\mathrm{d}\tau\mathop{}\!\mathrm{d}r$
$\leqslant C\gamma_{\max}\int_{0}^{t}c_{r}\mathbb{E}[\mathcal{E}_{r}]\gamma_{r}\mathcal{K}(t-r)\mathop{}\!\mathrm{d}r$
$=C\gamma_{\max}B(t),$
where the last inequality follows from Lemma C.9. Therefore, we get
$B(t)$ $\leqslant\frac{1}{1-C\gamma_{\max}}\int_{0}^{t}\mathcal{K}(t-\tau)\,c_{\tau}\gamma_{\tau}\left(g_{s}(\tau)+M^{-s\beta}+\sigma^{2}\int_{0}^{\tau}\mathcal{K}(\tau-r)c_{r}\gamma_{r}\mathop{}\!\mathrm{d}r\right)\mathop{}\!\mathrm{d}\tau$
$B(t)$ $\geqslant\int_{0}^{t}\mathcal{K}(t-\tau)\,c_{\tau}\gamma_{\tau}\left(g_{s}(\tau)+M^{-s\beta}+\sigma^{2}\int_{0}^{\tau}\mathcal{K}(\tau-r)c_{r}\gamma_{r}\mathop{}\!\mathrm{d}r\right)\mathop{}\!\mathrm{d}\tau.$
As $C\gamma_{\max}\leqslant\frac{1}{2}$, we have
$B(t)\eqsim\int_{0}^{t}\mathcal{K}(t-\tau)\,c_{\tau}\gamma_{\tau}\left(g_{s}(\tau)+M^{-s\beta}+\sigma^{2}\int_{0}^{\tau}\mathcal{K}(\tau-r)c_{r}\gamma_{r}\mathop{}\!\mathrm{d}r\right)\mathop{}\!\mathrm{d}\tau.$
Let
$p(t):=\int_{0}^{t}g_{2-1/\beta}(\tau)\gamma_{\tau}\mathop{}\!\mathrm{d}\tau,\quad q(t):=\int_{0}^{t}g_{2-1/\beta}(t-\tau)g_{s}(\tau)\gamma_{\tau}\mathop{}\!\mathrm{d}\tau.$ (22)
Noting $\mathcal{K}(t)=g_{2-1/\beta}(t)$, we have

### v3

@@ C.3 The Case of Top-$M$ Features
It holds for $s\leqslant 1-\frac{1}{\beta}$ that
$\int_{0}^{t}(\mathcal{K}_{\infty}(t-\tau)-\mathcal{K}_{M}(t-\tau))\gamma_{\tau}\mathop{}\!\mathrm{d}\tau\lesssim M^{-\beta+1}.$ (41)
Proof of Lemma.
First from the definition of $\mathcal{K}$ we obtain
$\mathcal{K}_{\infty}(t)-\mathcal{K}_{M}(t)=\sum_{j=M+1}^{\infty}\lambda_{j}^{2}e^{-2\lambda_{j}t}.$
Therefore, we have
$\int_{0}^{t}(\mathcal{K}_{\infty}(t-\tau)-\mathcal{K}_{M}(t-\tau))\gamma_{\tau}\mathop{}\!\mathrm{d}\tau$
$=\gamma_{\max}\int_{0}^{\infty}\sum_{j=1}^{M}\lambda_{j}^{2}e^{-2\lambda_{j}\tau}\mathop{}\!\mathrm{d}\tau$
$\eqsim\gamma_{\max}\sum_{j=1}^{M}j^{-\beta}\lesssim\gamma_{\max}M^{-\beta+1}.$
By $s\leqslant 1-\frac{1}{\beta}$ and that $\gamma_{\max}$ is sufficiently small, we can combine the upper bound with (41), (40), and directly conclude that
@@
$\mathbb{E}[\mathcal{E}_{t}]$ $\lesssim M^{-s\beta}+e_{M}(t)+\int_{0}^{t}\mathcal{K}_{\infty}(t-\tau)(e_{M}(\tau)+\sigma^{2})\gamma_{\tau}\mathop{}\!\mathrm{d}\tau$
$\lesssim M^{-s\beta}+e_{M}(t)+\int_{0}^{t}\mathcal{K}_{M}(t-\tau)(e_{M}(\tau)+\sigma^{2})\gamma_{\tau}\mathop{}\!\mathrm{d}\tau$
$\qquad+\int_{0}^{t}(\mathcal{K}_{\infty}(t-\tau)-\mathcal{K}_{M}(t-\tau))(e_{M}(\tau)+\sigma^{2})\gamma_{\tau}\mathop{}\!\mathrm{d}\tau$
[…]
$\qquad+M^{\max\{-s\beta,-2\beta+1\}}+\sigma^{2}M^{-\beta+1}$
$\lesssim M^{-s\beta}+e_{M}(t)+\int_{0}^{t}\mathcal{K}_{M}(t-\tau)(e_{M}(\tau)+\sigma^{2})\gamma_{\tau}\mathop{}\!\mathrm{d}\tau.$
Combining with the lower bound in FSL Theorem 4.3, the result follows. ∎
C.4 The Case of Random-$M$ Features
Notations. Throughout this section, for a power series $P(x)=\sum_{i=0}^{\infty}a_{i}x^{i},$ we write $P(\mathbf{A})=\sum_{i=0}^{\infty}a_{i}\mathbf{A}^{i}$ to denote the result of substituting a square matrix $\mathbf{A}\in\mathbb{R}^{d\times d}$ into $P(\cdot)$, where the power $\mathbf{A}^{i}$ is obtained through matrix multiplications.
For a vector $\mathbf{v}\in\mathbb{R}^{d}$, we write the norm $\lVert\mathbf{v}\rVert$ for $\lVert\mathbf{v}\rVert_{2}$ if not otherwise specified. For a matrix $\mathbf{A}\in\mathbb{R}^{d\times d}$, we use $\lVert\mathbf{A}\rVert_{2}$ to denote the operator norm induced by the vector 2-norm, and $\mu_{1}(\mathbf{A})\geqslant\mu_{2}(\mathbf{A})\geqslant\dots\geqslant\mu_{d}(\mathbf{A})$ to denote the eigenvalues of $\mathbf{A}$. For a positive semi-definite matrix $\mathbf{C}\in\mathbb{R}^{d\times d}$ and a vector $\mathbf{v}\in\mathbb{R}^{d}$, we define $\lVert\mathbf{v}\rVert_{\mathbf{C}}$ as
$\lVert\mathbf{v}\rVert_{\mathbf{C}}=\mathbf{v}^{\top}\mathbf{C}\mathbf{v}.$
First, for the random-$M$ feature setting, we can establish the following Volterra equation.
Proposition C.11.
Suppose $0<s\leqslant 1$. Then, with probability at least $1-e^{-\Omega(M)}$, a similar Volterra equation as derived in Theorem C.6 continues to hold for the random-feature case; that is,